What are the equilibrium concentrations of H2SO3, H+, HSO3−, and SO32−in a 0.050 M solution of sulfurous acid H2SO3 at 25 oC? For H2SO3 at 25 oC, Ka1 = 1.5×10−2 and Ka2 = 1.0×10−7
1
Expert's answer
2013-02-28T10:16:46-0500
Concentration of H2SO3 is 0.050 M
Concentration of HSO3- is next:
K1=[H+][HSO3-]/[H2SO4][H+]=[HSO3-]
K1=[HSO3-]^2/[H2SO4]
[HSO3-]^2=K1*[H2SO4]
[HSO3-]=(K1*[H2SO4])^1/2
[HSO3-]=(0.015*0.05)^1/2=2.77*10^-2
The same with [SO42-]
[SO32-]=(K2*[SO42-])^1/2 = 5.26*10^-5
[H+] is the sum of this two results